2x-5(x-10)is less than and = to x+3
2x-5x-50 < and equal to x+3
-3x-50< and equal to x+3
I get to here and I don't know what to do with all the x's and what not. can someone help me?

Answer by checkley77(12844)

2x-5(x-10)is less than and = to x+3
2X-5X+50<=X+3
-3X-X<=3-50
-4X<=-47
X>=-47/-4 (NOTE : WHEN YOU DIVIDE BY A NEGATRIVE VALUE THE < SIGN CHANGES TO A > SIGN.
X>=11.75 ANSWER.
PROOF:
LET X=12.
2*12-5(12-10)<=12+3
24-5(2)<=15
24-10<=15
14<=15